Mason Company has two manufacturing departments—Machining and Assembly. The company considers all of its manufacturing overhead costs to be fixed costs. It provided the following estimates at the beginning of the year as well as the following information with respect to Jobs A and B: Estimated Data Machining Assembly Total Manufacturing overhead $ 9,522,000 $ 529,000 $ 10,051,000 Direct labor-hours 23,000 414,000 437,000 Machine-hours 414,000 18,000 432,000 Job A Machining Assembly Total Direct labor-hours 5 10 15 Machine-hours 11 2 13 Job B Machining Assembly Total Direct labor-hours 4 5 9 Machine-hours 12 3 15 Required: 1. If Mason Company uses a plantwide predetermined overhead rate with direct labor-hours as the allocation base, how much manufacturing overhead cost would be applied to Job A? Job B? 2. Assume that Mason Company uses departmental predetermined overhead rates. The Machining Department is allocated based on machine-hours and the Assembly Department is allocated based on direct labor-hours. How much manufacturing overhead cost would be applied to Job A? Job B? (Round your intermediate calculations and final answers to 2 decimal places.)
Answer #1

1) Plantwide overhead rate = 10051000/437000 = 23 per DLH

Manufacturing overhead applied to Job A = 15*23 = 345

Manufacturing overhead applied to Job B = 9*23 = 207

2) Machining overhead rate = 9522000/414000 = 23 per MH

Assembly overhead rate = 529000/414000 = 1.28 per DLH

Job A = 11*23+10*1.28 = 265.80 or 266

Job B = 12*23+5*1.28 = 282.40 or 282